### v1

# Learning Manipulation Skills through Robot Chain-of-Thought with Sparse Failure Guidance

Abstract
The acquisition of manipulation skills through language instruction remains an unresolved challenge. Recently, vision-language models have made significant progress in teaching robots these skills. However, their performance is restricted to a narrow range of simple tasks. In this paper, we propose that vision-language models can provide a superior source of rewards for agents. Our method decomposes complex tasks into simpler sub-goals, enabling better task comprehension and avoiding potential failures with sparse failure guidance. Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP. Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rate compared to the best baseline, RoboCLIP, across a series of manipulation tasks. It has shown a comprehensive understanding of a wide range of robotic manipulation tasks.
1 Introduction
Learning manipulation skills from language instruction is a critical challenge in robot learning. Despite advancements (RT1; RT2; RTH), mastering robot manipulation remains an unresolved problem. Recently, vision-language models such as CLIP (CLIP), LIV (LIV), Flamingo (Flamingo), PaLM (PaLM), and GPT-4V have markedly improved robots’ capabilities in acquiring manipulation skills (LIV; RoboFlamingo; PaLME). However, these models are limited to a narrow range of simple tasks, such as closing a drawer. Additionally, CLIP (MetaCLIP) and LIV (LIV) focus on language and visual goal alignment but lack an understanding of the manipulation process. Conversely, RoboCLIP (RoboCLIP) provides a sparse reward signal using S3D (S3D) which lacks spatial reasoning capabilities and task completion awareness.
This study proposes a superior reward design mechanism that leverages pre-trained vision-language models as a source of rewards. Vision language foundation models can provide prior knowledge for robots on how to manipulate objects and reinforce the task completion process. Consequently, our algorithm comprehends a wide range of manipulation tasks.
By analogy, consider how children learn. They break complex tasks into manageable stages, simplifying the task and facilitating its completion. For instance, to open a door, children decompose the task into sub-goals such as approaching the handle, grasping it, pulling it back, and opening the door. During learning, they also draw lessons from potential failures and successes, guided by parental advice and reinforcement.
Our algorithm draws inspiration from child learning paradigms. To better understand the manipulation task process, we use a robot chain-of-thought to decompose tasks into a series of actionable steps, leveraging prior task knowledge from foundation models. Meanwhile, we provide sparse human failure guidance, highlighting potential failures for the agents. The designed reward mechanism yields higher rewards when the robot performs tasks in line with task descriptions while avoiding potential failure cases. To reinforce the task completion process, we collect success experiences through foundation models to strengthen the policy.
In this paper, we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RTFG). Our RTFG introduces an innovative reward feedback mechanism using vision language foundation models to equip robots with a broad spectrum of manipulation skills. We demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline. Ablation studies show that each component of our algorithm contributes to the overall performance.
2 Related Work
2.1 Learning Manipulation Skills for Robotics
Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) (perceiver; CLIPort; octo) and inverse reinforcement learning (IRL) (GAIL; liu_imitation). BC directly maps the state space to the action space to mimic the expert’s actions. However, BC faces several limitations, including compounding errors (DAgger), the copycat problem (copycat), and out-of-distribution sample shifts (fight_fire). In contrast, IRL derives both the reward function and the policy from expert demonstrations. However, IRL struggles with sample efficiency, particularly in real-world applications (jeong2020self; matas2018sim; xu2023roboninja).
Mastering manipulation skills with language inputs directly presents several challenges, such as a lack of reward signals and inefficiency in understanding vision-language contexts. Recent studies have explored integrating foundation models to enhance the understanding of language instructions (SayCan). Our algorithm is designed to learn manipulation skills from language inputs by effectively leveraging pre-trained vision-language models as a source of rewards.
2.2 Foundation Models for Robotics
Vision-language foundation models equip robot agents with compact pre-trained representations. R3M (R3M) co-trains video and language instructions in a self-supervised manner (TCN), enhancing the robot’s ability to mimic expert actions. PaLM-E (PaLME) integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process. RT-1 (RT1) uses transformers to better mimic expert behaviors. RoboFlamingo (RoboFlamingo) uses multi-step observation inputs to capture a comprehensive view of observational data. RT-2 (RT2) innovates by conceptualizing robot actions as a form of language. Furthermore, RT-H (RTH) introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large pre-trained language models. However, while capable of high-level task decomposition, prior models struggle with learning low-level action primitives. Our algorithm addresses this by employing reinforcement learning to master low-level action primitives. Additionally, prior models often lack mechanisms for correcting potential failure cases. Our algorithm addresses this by incorporating sparse failure guidance, thereby strengthening the robustness of our method.
Foundation models significantly enhance robot manipulation capabilities by offering valuable reward signals. CLIP (CLIP), for instance, aligns image representations with language embeddings. Consequently, (MetaCLIP) employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal. LIV (LIV) innovatively combines VIP (VIP) and CLIP to more effectively learn both visual and linguistic goal representations. RoboCLIP (RoboCLIP) adopts the S3D (S3D) representation model to assess the reward for the final observation of a trajectory. Furthermore, RL-VLM-F (RL_VLM_F) directly utilizes GPT-4V to label preferences for agent rollout data.
However, current models still struggle to master a broad spectrum of robot tasks, often succeeding only in simple tasks such as closing a drawer. Vision-language models like CLIP (CLIP) and LIV (LIV) lack an understanding of the manipulation process. While RoboCLIP (RoboCLIP) offers a sparse reward signal that aids in policy learning, it lacks spatial reasoning capabilities and task completion awareness. RL-VLM-F (RL_VLM_F) excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities. In contrast, our algorithm not only comprehends detailed language instructions from the robot chain-of-thought but also considers failure scenarios. Additionally, it leverages VLM’s spatial reasoning ability through success experience regularization, offering a more comprehensive approach to mastering robot tasks.
Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using the VLM’s chain-of-thought processing, breaking it down into several detailed prompts. The robot also considers some failure guidance. Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills. Successful experiences are recorded to reinforce effective behaviors using VLM.
3 Background
We consider scenarios where the robotic system receives specific verbal commands, such as "open the door." To model this learning process, we employ a Markov Decision Process (MDP) defined by $(S,A,r,P,\gamma)$, where $S$ denotes the state space, $A$ the action space, $r$ the reward model, $P$ the transition model, and $\gamma$ the discount factor.
Chain-of-Thought (CoT) The CoT (CoT) approach is a reasoning method designed to enhance the problem-solving capabilities of large language models. This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving. By explicitly detailing these intermediate steps, large language models such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
VideoCLIP VideoCLIP (VideoCLIP) is a multi-modal model that connects visual and textual information, specifically designed to understand videos by associating them with relevant text descriptions. The model is trained on a diverse dataset of videos and associated texts, allowing it to develop a nuanced understanding of the relationship between video segments and language. In particular, VideoCLIP minimizes the sum of two multi-modal contrastive losses:
$\mathcal{L}=-\sum_{(v,t)\in\mathcal{B}}(\log\text{NCE}(z_{v},z_{t})+\log\text{NCE}(z_{t},z_{v})),$ (1)
where $\mathcal{B}$ is the batch containing sampled video-text pairs, and NCE$(z_{v},z_{t})$ and NCE$(z_{t},z_{v})$ correspond to the contrastive loss on video-to-text similarity and vice versa. Specifically, the video-to-text contrastive loss is given by:
$\text{NCE}(z_{v},z_{t})=\frac{\exp(z_{v}\cdot z_{t}^{+}/\tau)}{\sum_{z\in{\{z_{t}^{+},z_{t}^{-}\}}}\exp(z_{v}\cdot z/\tau)},$ (2)
where $\tau$ is a temperature hyper-parameter, $z_{t}^{+}$ and $z_{t}^{-}$ are positive and negative text embeddings respectively, and $z_{v}$ is a video segment embedding.
4 Method
Figure 2: First, text embeddings are generated using VLM chain-of-thought and human failure guidance. Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings. Finally, VLM evaluates the final observation. If this observation is deemed successful, this trajectory is recorded in the success experience replay buffer for policy regularization.
In section 4.1, we introduce the robot chain-of-thought interpretation of the command and sparse human failure guidance. In section 4.2, we explain how to define the rewards for policy learning with foundation models. Next, in section 4.3, we discuss policy learning with success experience regularization. Finally, section 4.4 outlines the pipeline of our algorithm.
4.1 Robot Chain-of-Thought with Sparse Failure Guidance
Manipulation skills consist of a series of actionable control steps. For instance, opening a door includes sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door. Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control. However, robots may still encounter potential failures during learning. Offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $z_{l}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
In our algorithm, $z_{l}^{-}$ is a set of offline potential failure case descriptions consisting of 2 to 4 prompts. To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training the agent.
4.2 Foundation Models as Rewarder
The agent acquires manipulation skills by interacting with an environment, guided by language inputs. A central challenge we address is defining a reward function that effectively reinforces desired behaviors. In this work, we argue that pre-trained vision-language models can provide the agent with prior task knowledge, such as how to manipulate objects. Consequently, we propose a novel reward design mechanism based on robot CoT and sparse failure guidance. Specifically, we explore the usage of VideoCLIP (VideoCLIP) as a vision-language foundation model. VideoCLIP excels in capturing the temporal dynamics of video segments and interpreting associated language prompts, aiming to align robot movements with actions described in the task descriptions.
The reward signal for learning manipulation skills is defined as follows:
$\text{reward}=\frac{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)}{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (3)
where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task descriptions from robot CoT interpretation, and $z_{l}^{-}$ is a set of potential failure case descriptions guided by humans. The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims to distinguish between video segments and corresponding positive or negative texts. We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP (RoboCLIP). Moreover, the max operation in Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent. Otherwise, the reward signal would penalize undesired behaviors.
4.3 Policy Learning with Success Experience Regularization
Our reward design mechanism motivates the agent to explore the necessary manipulation skills. Although this reward signal performs well in manipulation tasks, the VideoCLIP rewards do not reflect the completion of the manipulation tasks. They only indicate whether the robot’s movements align with detailed language goals. To better reinforce task completion, we introduce policy learning with success experience regularization.
We employ GPT-4V to assess the success of the agent’s final observation in its rollout trajectory. If it deemed successful, this trajectory is archived into a success experience replay buffer, enabling the agent to reinforce this success. Subsequently, we introduce a regularization mechanism aimed at refining the agent’s policy to reinforce successful behavior. The regularization loss is integrated with the policy loss, formulated as follows:
=RL $\mathcal{L}=\text{RL Loss}+\lambda\sum_{(s,a)\in D_{E}}-\log(\pi(a|s)),$ (4)
where $D_{E}$ is the success experience replay buffer and $\lambda$ is the hyper-parameter. By incorporating success experiences into the policy loss calculation, the agent is encouraged to repeat desired behaviors. However, the reward signal for success experiences should also be reinforced to help the agent understand the task completion signal. Therefore, we correct the reward function as follows: $r(s_{T})=r(s_{T})+100$, where $s_{T}$ is the successful final observation of the rollout trajectory.
4.4 Pipeline
Our algorithm begins with GPT-4V interpreting language instructions into detailed task descriptions. Meanwhile, humans guide the agent with several potential failure case descriptions. Subsequently, rewards are calculated following Equation 3 with a moving window that slides over the agent’s rollout trajectories. The agent is then updated with these learned rewards to further develop the skills. Concurrently, GPT-4V evaluates the final observation of the rollout trajectory and assigns a preference label, indicating whether it is a successful observation. If it is successful, we record the trajectory into a success experience replay buffer for policy regularization and correct the reward function.
Figure 2 illustrates our method for computing rewards and success regularization that aid the robot in learning manipulation skills. This involves a moving window that traverses the agent’s rollout trajectory, the computation of NCE rewards using language prompts, and the policy regularization using success experiences gathered through GPT-4V. This comprehensive approach aims to effectively teach the agent desired manipulation skills through reinforcement learning.
5 Experiment
Figure 3: Training curves for baselines and our algorithm within 3 random seeds. Each data point is evaluated with 20 sampled trajectories. The shaded area displays the range of one standard deviation. Our algorithm gets best performance compared to other baselines, achieves 5.4×5.4$\times$5.4 × improvement compared to RoboCLIP.
In this section, we assess the effectiveness of our proposed algorithm by conducting experiments in comparison with several baseline models. The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms? The results are presented in section 5.1. (2) What are the contributing factors in our algorithm? The ablation studies are presented in section 5.2.
Environments To evaluate the effectiveness of our proposed algorithm, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment. We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by (ROT). Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 (DrQv2) as the reinforcement learning algorithm. Detailed implementation specifics can be found in the references: CLIP (MetaCLIP), LIV (LIV), and RoboCLIP (RoboCLIP). All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs. Implementation details are provided in the Appendix.
5.1 Overall Comparison
To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP (MetaCLIP), LIV (LIV), and RoboCLIP (RoboCLIP). CLIP (as described in (MetaCLIP)) employs CLIP (CLIP) as the vision-language model to measure the distance between current observations and language goals. LIV (LIV) aligns goal representations with language instructions, and RoboCLIP (RoboCLIP) utilizes S3D (S3D) to measure the distance between trajectory videos and language as a form of reward. However, CLIP (CLIP) and LIV (LIV) lack an understanding of the manipulation process, while RoboCLIP (RoboCLIP) lacks spatial reasoning capabilities and task completion awareness. Figure 3 illustrates the training curves, depicting the success rates for each algorithm. Each evaluation involves 20 trajectories, and the experiments were replicated across three different environmental random seeds to introduce variability in the position of objects.
Our findings indicate distinct performance trends across various tasks. RoboCLIP displays promising results in tasks involving drawer and window manipulation, as well as soccer, but it is most effective in the simplest tasks such as drawer closing. CLIP performs well in tasks related to faucet, drawer and window manipulation, and soccer, indicating that CLIP focuses more on task completion rather than the manipulation process. LIV shows competence primarily in tasks involving drawer closure, window opening, and soccer. Our algorithm consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors. These results underscore the effectiveness of our approach in handling diverse and challenging scenarios.
5.2 Ablation Study
In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and success regularization). The results indicate that the most contributing factor to our algorithm’s performance is success regularization, followed by human failure guidance. Each component contributes to the overall performance of our algorithm. Additionally, we conduct ablation studies comparing GPT-4V success labels with ground truth success labels, demonstrating the robustness of our GPT-4V labeling. Furthermore, we explore different reward design mechanisms, showing that they can also achieve high performance within our algorithm. Finally, experiments with all baselines incorporating success regularization reveal that our algorithm’s superiority is not solely due to success regularization but also the reward design mechanism.
Question 1: how significant is each component of our algorithm in contributing to the results?
Figure 4: Final policy performance for ablations and our algorithm across 3 random seeds. The results show that success regularization is the most significant contributing factor in our algorithm, followed by human failure guidance. Each component contributes to the overall performance of our algorithm.
Our algorithm incorporates several critical components: robot chain-of-thought processing via GPT-4V, sparse human failure guidance, and success experience regularization. To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance. A decline in performance upon the removal of any component demonstrates its essential role in our algorithm.
Figure 4 illustrates the success rates of the final policy performance of our algorithm compared to its ablated versions. The labels for these ablations are: w/oReg (without success experience regularization), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
The results confirm the critical importance of each component in our algorithm, with success experience regularization being particularly pivotal. Human failure guidance is the second most important component, motivating the agent to consider potential failure cases. Additionally, robot chain-of-thought processing helps the agent better understand the process of manipulation tasks, further contributing to the performance of our algorithm.
Question 2: what is the empirical difference between the labels assigned by GPT-4V and the ground truth labels in the context of success experience collection?
Figure 5: Ablation study for our algorithm with different success labeling methods (GPT-4V labeling or ground truth (GT) labeling). The results demonstrate that the success labels generated by GPT-4V are robust.
Figure 6: Ablation study for our algorithm with CoT as one single prompt or multiple prompts. The results show that single prompt for reward mechanism also works well.
Our algorithm incorporates GPT-4V to gather successful trajectories for policy regularization. To validate the robustness of our GPT-4V labeling in collecting success experiences, we conduct experiments comparing GPT-4V labels and ground truth labels. If the performance using GPT-4V labels is similar to that using ground truth labels, it would confirm that GPT-4V is robust in collecting success experiences.
The process for GPT-4V analysis is outlined in Figure 2. Initially, we present GPT-4V with a scenario indicating a successful state (e.g., the door is open). We then provide a prompt: "Consider the following image; this is a simulation environment. In this image, the door of the box is open. We define it when the door is open, the output is 1. Otherwise, the output is 0." Subsequently, we present the final observation to GPT-4V and prompt, "What is the output of this image?" We collect the trajectory as successful when GPT-4V’s response is 1.
Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labeling methods for success experience collection. Our findings show that the GPT-4V labeling method achieves performance comparable to the ground truth across various tasks, demonstrating the robustness of our GPT-4V success labels.
Question 3: what is the empirical difference between our algorithm and the one with a single chain-of-thought interpreted positive prompt?
Our algorithm computes the NCE-rewards based on the maximum distance between the moving window video and multiple positive prompts. This measurement indicates how far the moving window video is from the language goal derived from robot chain-of-thought interpretations. Conversely, our ablation method computes the NCE-rewards based on the distance between the moving window video and a single interpreted positive prompt. The NCE-reward function is shown as:
$\text{reward}=\frac{\exp(z_{v}\cdot z_{l}^{+}/\tau)}{\exp(z_{v}\cdot z_{l}^{+}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (5)
where $z_{l}^{+}$ indicates the single positive prompt from the GPT-4V interpretation. For example, the prompt for opening a door is "the robot hand first approaches the door’s handle, then grasps it, pulls it back, and finally opens the door."
Figure 6 shows the comparison between our algorithm with multiple positive prompts and the single positive prompt ablation. The empirical results demonstrate that our algorithm achieves comparable results with a single positive prompt. It indicates that our algorithm is robust to the setting of multiple or single positive prompts.
Figure 7: Ablation study for all baselines with success experience regularization. The results show that our algorithm performs well not only because of the success regularization component but also due to the reward design mechanism.
Question 4: what is the empirical difference between our algorithm and all baselines with success experience regularization?
Question 1 shows that success experience regularization component plays Question 1 shows that the success experience regularization component plays the most important role in our algorithm. To evaluate the significance of the reward design mechanism, we conduct experiments applying success regularization to all baselines. If the results show that our algorithm still outperforms other baselines with success regularization, then it proves that our algorithm’s superiority is due not only to the success regularization but also to the reward design mechanism.
Figure 7 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include success experience regularization. The results demonstrate that our algorithm outperforms other baselines on almost all tasks. This indicates that our algorithm’s superiority is due not only to the success regularization component but also to the reward design mechanism.
6 Conclusion
In this study, we present a novel algorithm for robotic manipulation skill acquisition. We propose that pre-trained vision-language models can instruct robots on prior knowledge essential for manipulation tasks. Our algorithm simplifies complex tasks and acquires manipulation skills through a robot’s chain-of-thought process with sparse failure guidance, mirroring the developmental learning trajectory of a child. Specifically, our algorithm is designed to extract intricate temporal visual cues from the robot’s thought process, enabling a nuanced understanding of the task at hand. It also incorporates a learning strategy by describing potential failure cases through sparse failure guidance, which enhances its manipulation policies. Through success experience regularization, our algorithm strengthens its policies, thereby improving performance. Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage over established baseline models. This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic manipulation tasks in the future.
7 Appendix
7.1 Environment Setup
We use MetaWorld-v2 as our test robotic environments. We choose 10 tasks as by [ROT] to evaluate our algorithm and baselines. They are drawer close, window close, door close, faucet close, button press, drawer open, window open, door open, faucet open and soccer. Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
To reduce the exploration space, we set the gripper a new position. With varying random seeds, the objects like drawer, window, door, faucet, button, and soccer would vary its positions. The position coordinate are provided in Table 1.
Table 1: Position coordinates for robot gripper as setting with MetaWorld-v2.
Task Coordinate
Drawer Close $0,0.6,0.2$
Drawer Open $0,0.8,0.05$
Window Close $0.25,0.80,0.16$
Window Open $-0.15,0.77,0.16$
Door Close $-0.5,0.6,0.2$
Door Open $0,0.6,0.2$
Faucet Close 0.,0.6,0.10.,0.6,0.10 . , 0.6 , 0.1
Faucet Open 0.,0.6,0.10.,0.6,0.10 . , 0.6 , 0.1
Button Press $0,0.6,0.2$
Soccer $0.6,0$ 0.6 , 0.06
Since the default camera view is not suitable for our algorithms. We re-set the camera for training all the experiments. The principle for choosing camera is that we set the object and gripper in the view. The camera coordinates are provided in Table 2.
Table 2: Position coordinates and angles for camera as setting with MetaWorld-v2.
Task Coordinate Angle
Door Close $-1.1,-0.4,0.6$
Other Tasks $0.55,0.3,0.5$ $3.9,2.3,0.6$
7.2 GPT-4V Success Labeling Details
Figure 8: The successful trajectories are gathered through GPT-4V success labeling. GPT-4V would label the success preference to the final observation of agent’s rollout trajectory. If it deemed successful, the trajectory would be recorded into a replay buffer.
Our algorithm incorporates GPT-4V to gather successful trajectories for policy regularization. The process for GPT-4V analysis is outlined in Figure 8. Initially, we present GPT-4V with a scenario indicating a successful state (e.g., the door is open). We then provide a prompt: "Consider the following image; this is a simulation environment. In this image, the door of the box is open. We define it when the door is open, the output is 1. Otherwise, the output is 0." Subsequently, we present the final observation to GPT-4V and prompt, "What is the output of this image?" We collect the trajectory as successful when GPT-4V’s response is 1.
To reduce the query times with GPT-4V, we distill a classifier from GPT-4V labeling data. First, we use GPT-4V to label 300 pictures with diverse final observations. Next, we fine-tune a ResNet-18 classifier based on this dataset. Before training, we use some augmentation techniques such as random crop resize to augment the dataset. Finally, we use this classifier to classify the final observation as successful or not for instance.
7.3 Task Robot Chain-of-Thought Prompt
Manipulation skills consist of a series of actionable control steps. For instance, opening a door includes sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door. Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control.
We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $z_{l}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks. The task robot CoT prompts are provided in Table 3.
Table 3: Prompts from robot chain-of-thought for task interpretation.
Task Robot Chain-of-Thought Prompts
Drawer Close the robot hand pushes the drawer;
the drawer is closed.
Drawer Open the robot hand moves to the top of the drawer’s handle;
the robot hand moves inside the drawer’s handle;
the robot hand pulls back the drawer’s handle;
the robot hand opens the drawer.
Window Close the robot hand moves to the left of the window’s handle;
the robot hand pushes the handle towards right;
the robot hand closes the window.
Window Open the robot hand moves to the right of the window’s handle;
the robot hand pushes the handle towards left;
the robot hand opens the window.
Door Close the robot hand pushes the door close;
the door is closed.
Door Open the robot hand is at the front of the door;
the robot hand moves to the left of the door’s handle;
the robot hand pulls back the door’s handle;
the robot hand opens the door
Faucet Close the robot hand moves to the left of the faucet;
the robot hand pulls the faucet towards right;
the faucet is closed.
Faucet Open the robot hand moves to the right of the faucet;
the robot hand pulls the faucet towards left;
the faucet is open.
Button Press the robot hand moves towards the red button;
the robot hand pushes the button;
the button is pressed.
Soccer the robot hand moves towards soccer;
the robot hand pushes soccer into the net;
the soccer is in the net.
7.4 Sparse Human Failure Guidance
Robots may encounter potential failures during learning. Offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
$z_{l}^{-}$ is a set of offline potential failure case descriptions consisting of 2 to 4 prompts. To enable the agent to recognize and adjust to potential failures, we provide sparse human failure guidance before training the agent. The sparse human failure guidance are provided in Table 4.
Table 4: Sparse human failure guidance.
Task Human Failure Guidance
Drawer Close the robot hand is moving away from the drawer;
the robot hand is at the left of the drawer;
the robot hand is at the right of the drawer.
Drawer Open the robot hand is at the right of the drawer’s handle;
the robot hand is at the left of the drawer’s handle;
the robot hand is under the drawer’s handle.
Window Close the robot hand is at the top of the window’s handle;
the robot hand is at the right of the window’s handle;
the robot hand is under the window’s handle.
Window Open the robot hand is at the top of the window’s handle;
the robot hand is at the left of the window’s handle;
the robot hand is under the window’s handle.
Door Close the robot hand is moving away from the door;
the robot hand is at the left of the door.
Door Open the robot hand is at the right of the door’s handle;
the robot hand is at the top of the door’s handle;
the robot hand is under the door’s handle;
the robot hand is at the left of the box.
Faucet Close the robot hand is to the right of the faucet;
the robot hand is at the top of the faucet.
Faucet Open the robot hand is to the left of the faucet;
the robot hand is at the top of the faucet.
Button Press the robot hand moves away from the red button;
the robot hand is at the left of the red button;
the robot hand is at the right of the red button.
Soccer the robot hand is moving away from soccer;
the robot hand is at the top of soccer.

### v2

@@
 Abstract
-The acquisition of manipulation skills through language instruction remains an unresolved challenge.
-Recently, vision-language models have made significant progress in teaching robots these skills.
-However, their performance is restricted to a narrow range of simple tasks.
-In this paper, we propose that vision-language models can provide a superior source of rewards for agents.
-Our method decomposes complex tasks into simpler sub-goals, enabling better task comprehension and avoiding potential failures with sparse failure guidance.
+Defining reward functions for skill learning has been a long-standing challenge in robotics.
+Recently, vision-language models (VLMs) have shown promise in defining reward signals for teaching robots manipulation skills.
+However, existing works often provide reward guidance that is too coarse, leading to inefficient learning processes.
+In this paper, we address this issue by implementing more fine-grained reward guidance.
+We decompose tasks into simpler sub-tasks, using this decomposition to offer more informative reward guidance with VLMs.
+We also propose a VLM-based self imitation learning process to speed up learning.
 Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP.
 Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rate compared to the best baseline, RoboCLIP, across a series of manipulation tasks.
-It has shown a comprehensive understanding of a wide range of robotic manipulation tasks.
+Keywords: Manipulation Skills, Vision Language Models, Chain-of-Thought
 1 Introduction
-Learning manipulation skills from language instruction is a critical challenge in robot learning.
-Despite advancements (RT1; RT2; RTH), mastering robot manipulation remains an unresolved problem.
-Recently, vision-language models such as CLIP (CLIP), LIV (LIV), Flamingo (Flamingo), PaLM (PaLM), and GPT-4V have markedly improved robots’ capabilities in acquiring manipulation skills (LIV; RoboFlamingo; PaLME).
-However, these models are limited to a narrow range of simple tasks, such as closing a drawer.
-Additionally, CLIP (MetaCLIP) and LIV (LIV) focus on language and visual goal alignment but lack an understanding of the manipulation process.
-Conversely, RoboCLIP (RoboCLIP) provides a sparse reward signal using S3D (S3D) which lacks spatial reasoning capabilities and task completion awareness.
-This study proposes a superior reward design mechanism that leverages pre-trained vision-language models as a source of rewards.
-Vision language foundation models can provide prior knowledge for robots on how to manipulate objects and reinforce the task completion process.
-Consequently, our algorithm comprehends a wide range of manipulation tasks.
-By analogy, consider how children learn.
-They break complex tasks into manageable stages, simplifying the task and facilitating its completion.
-For instance, to open a door, children decompose the task into sub-goals such as approaching the handle, grasping it, pulling it back, and opening the door.
-During learning, they also draw lessons from potential failures and successes, guided by parental advice and reinforcement.
-Our algorithm draws inspiration from child learning paradigms.
-To better understand the manipulation task process, we use a robot chain-of-thought to decompose tasks into a series of actionable steps, leveraging prior task knowledge from foundation models.
-Meanwhile, we provide sparse human failure guidance, highlighting potential failures for the agents.
-The designed reward mechanism yields higher rewards when the robot performs tasks in line with task descriptions while avoiding potential failure cases.
-To reinforce the task completion process, we collect success experiences through foundation models to strengthen the policy.
-In this paper, we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RTFG).
-Our RTFG introduces an innovative reward feedback mechanism using vision language foundation models to equip robots with a broad spectrum of manipulation skills.
-We demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline.
+Defining reward functions for learning robotic skills has been a long standing challenge in robotics [1].
+Traditional methods typically define the rewards through human knowledge and require significant engineering efforts.
+Some methods also use expert demonstrations to define rewards [1].
+However, this procedure requires expert supervision and can be infeasible in complex scenarios [2] .
+Recently, the success of vision language models (VLM) has inspired researchers to leverage its power to define rewards for skill learning.
+For example, RoboCLIP [2] captures the temporal information from robot movements with a video backbone S3D [3].
+In spite of the preliminary success of VLM-based reward modeling, a downside of these existing works is that their language guidance is too coarse for learning.
+Consider the task of opening the door, existing works only use the language embedding of "opening the door" in reward definition to guide the learning process.
+Such guidance can be too sparse for learning: when the robot is not able to approach the door at the beginning of learning process, the rewarder will always provide a near-zero feedback since the door is not likely to be opened.
+In contrast, in human world we typically use more fine-grained guidance such as "move towards the door", "move the hand towards the handle" and "grab the handle" during the teaching process.
+Intuitively, using such decomposition can make the guidance more precise, and will lead to richer and denser feedback.
+Based on this intuition, in this paper we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RoboCoT) to address this problem.
+Our main insight is to decompose the task description into a series of actionable steps and then leverage VLM to define a fine-grained reward to guide skill learning, which can make the learning process more sample efficient and yield higher success rates.
+We also propose a VLM-based self-imitation procedure to further speed up the learning process.
+In the experiments, we demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline.
 Ablation studies show that each component of our algorithm contributes to the overall performance.
+Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using a robot chain-of-thought processing, breaking it down into several detailed prompts.
+The robot also considers some failure guidance.
+Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills.
+Successful experiences are recorded by foundation models to reinforce effective behaviors (i.e., VLM-based self-imitation).
 2 Related Work
 2.1 Learning Manipulation Skills for Robotics
-Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) (perceiver; CLIPort; octo) and inverse reinforcement learning (IRL) (GAIL; liu_imitation).
+Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) [4, 5, 6] or inverse reinforcement learning (IRL) [7, 8].
 BC directly maps the state space to the action space to mimic the expert’s actions.
-However, BC faces several limitations, including compounding errors (DAgger), the copycat problem (copycat), and out-of-distribution sample shifts (fight_fire).
+However, BC faces several limitations, including compounding errors [9], the copycat problem [10], and out-of-distribution sample shifts [11].
 In contrast, IRL derives both the reward function and the policy from expert demonstrations.
-However, IRL struggles with sample efficiency, particularly in real-world applications (jeong2020self; matas2018sim; xu2023roboninja).
+However, IRL struggles with sample efficiency, particularly in real-world applications [12, 13, 14].
 Mastering manipulation skills with language inputs directly presents several challenges, such as a lack of reward signals and inefficiency in understanding vision-language contexts.
-Recent studies have explored integrating foundation models to enhance the understanding of language instructions (SayCan).
-Our algorithm is designed to learn manipulation skills from language inputs by effectively leveraging pre-trained vision-language models as a source of rewards.
+Recent studies have explored integrating VLMs to enhance the understanding of language instructions [15].
+Our algorithm is designed by effectively leveraging VLMs as a source of rewards.
 2.2 Foundation Models for Robotics
-Vision-language foundation models equip robot agents with compact pre-trained representations.
-R3M (R3M) co-trains video and language instructions in a self-supervised manner (TCN), enhancing the robot’s ability to mimic expert actions.
-PaLM-E (PaLME) integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process.
-RT-1 (RT1) uses transformers to better mimic expert behaviors.
-RoboFlamingo (RoboFlamingo) uses multi-step observation inputs to capture a comprehensive view of observational data.
-RT-2 (RT2) innovates by conceptualizing robot actions as a form of language.
-Furthermore, RT-H (RTH) introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large pre-trained language models.
-However, while capable of high-level task decomposition, prior models struggle with learning low-level action primitives.
-Our algorithm addresses this by employing reinforcement learning to master low-level action primitives.
-Additionally, prior models often lack mechanisms for correcting potential failure cases.
-Our algorithm addresses this by incorporating sparse failure guidance, thereby strengthening the robustness of our method.
-Foundation models significantly enhance robot manipulation capabilities by offering valuable reward signals.
-CLIP (CLIP), for instance, aligns image representations with language embeddings.
-Consequently, (MetaCLIP) employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal.
-LIV (LIV) innovatively combines VIP (VIP) and CLIP to more effectively learn both visual and linguistic goal representations.
-RoboCLIP (RoboCLIP) adopts the S3D (S3D) representation model to assess the reward for the final observation of a trajectory.
-Furthermore, RL-VLM-F (RL_VLM_F) directly utilizes GPT-4V to label preferences for agent rollout data.
+VLMs equip robot agents with compact pre-trained representations.
+R3M [16] co-trains video and language instructions in a self-supervised manner [17], enhancing the robot’s ability to mimic expert actions.
+PaLM-E [18] integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process.
+RT-1 [19] uses transformers to better mimic expert behaviors.
+RoboFlamingo [20] uses multi-step observation inputs to capture a comprehensive view of observational data.
+RT-2 [21] innovates by conceptualizing robot actions as a form of language.
+Furthermore, RT-H [22] introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large language models.
+VLMs significantly enhance robot manipulation capabilities by offering valuable reward signals.
+CLIP [23], for instance, aligns image representations with language embeddings.
+Consequently, [24] employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal.
+LIV [25] innovatively combines VIP [26] and CLIP to more effectively learn both visual and linguistic goal representations.
+RoboCLIP [2] adopts the S3D [3] representation model to assess the reward for the final observation of a trajectory.
+Furthermore, RL-VLM-F [27] directly utilizes GPT-4V to label preferences for agent rollout data.
 However, current models still struggle to master a broad spectrum of robot tasks, often succeeding only in simple tasks such as closing a drawer.
-Vision-language models like CLIP (CLIP) and LIV (LIV) lack an understanding of the manipulation process.
-While RoboCLIP (RoboCLIP) offers a sparse reward signal that aids in policy learning, it lacks spatial reasoning capabilities and task completion awareness.
-RL-VLM-F (RL_VLM_F) excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities.
+VLMs like CLIP [23] and LIV [25] lack an understanding of the manipulation process.
+On the other hand, RoboCLIP [2] lacks spatial reasoning capabilities and task completion awareness.
+RL-VLM-F [27] excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities.
 In contrast, our algorithm not only comprehends detailed language instructions from the robot chain-of-thought but also considers failure scenarios.
-Additionally, it leverages VLM’s spatial reasoning ability through success experience regularization, offering a more comprehensive approach to mastering robot tasks.
-Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using the VLM’s chain-of-thought processing, breaking it down into several detailed prompts.
-The robot also considers some failure guidance.
-Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills.
-Successful experiences are recorded to reinforce effective behaviors using VLM.
+Additionally, it leverages VLM’s task completion reasoning ability through self imitation process, offering a more comprehensive approach to master robot tasks.
 3 Background
 We consider scenarios where the robotic system receives specific verbal commands, such as "open the door." To model this learning process, we employ a Markov Decision Process (MDP) defined by $(S,A,r,P,\gamma)$, where $S$ denotes the state space, $A$ the action space, $r$ the reward model, $P$ the transition model, and $\gamma$ the discount factor.
-Chain-of-Thought (CoT) The CoT (CoT) approach is a reasoning method designed to enhance the problem-solving capabilities of large language models.
+Chain-of-Thought (CoT) The CoT [28] approach is a reasoning method designed to enhance the problem-solving capabilities of large language models.
 This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving.
 By explicitly detailing these intermediate steps, large language models such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
-VideoCLIP VideoCLIP (VideoCLIP) is a multi-modal model that connects visual and textual information, specifically designed to understand videos by associating them with relevant text descriptions.
+VideoCLIP VideoCLIP [29] is a multi-modal model that connects visual and textual information, specifically designed to understand videos by associating them with relevant text descriptions.
 The model is trained on a diverse dataset of videos and associated texts, allowing it to develop a nuanced understanding of the relationship between video segments and language.
@@
 4 Method
-Figure 2: First, text embeddings are generated using VLM chain-of-thought and human failure guidance.
+Figure 2: First, text embeddings are generated using robot chain-of-thought (positive prompts) and human failure guidance (negative prompts).
 Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings.
-Finally, VLM evaluates the final observation.
-If this observation is deemed successful, this trajectory is recorded in the success experience replay buffer for policy regularization.
-In section 4.1, we introduce the robot chain-of-thought interpretation of the command and sparse human failure guidance.
-In section 4.2, we explain how to define the rewards for policy learning with foundation models.
-Next, in section 4.3, we discuss policy learning with success experience regularization.
-Finally, section 4.4 outlines the pipeline of our algorithm.
-4.1 Robot Chain-of-Thought with Sparse Failure Guidance
+Finally, VLMs evaluate the final observation.
+If this observation is deemed successful, this trajectory is recorded in the success replay buffer for self imitation.
+In section 4.1, we introduce the robot chain-of-thought interpretation.
+In section 4.2, we explain how to define the rewards for policy learning with VLMs.
+Finally, in section 4.3, we introduce VLM-based self imitation procedure via success experience regularization.
+4.1 Robot Chain-of-Thought
 Manipulation skills consist of a series of actionable control steps.
@@
 Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control.
-However, robots may still encounter potential failures during learning.
-Offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
-We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $z_{l}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
-In our algorithm, $z_{l}^{-}$ is a set of offline potential failure case descriptions consisting of 2 to 4 prompts.
-To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training the agent.
-4.2 Foundation Models as Rewarder
-The agent acquires manipulation skills by interacting with an environment, guided by language inputs.
+We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $l_{i}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
+4.2 VLMs as Rewarder
 A central challenge we address is defining a reward function that effectively reinforces desired behaviors.
-In this work, we argue that pre-trained vision-language models can provide the agent with prior task knowledge, such as how to manipulate objects.
-Consequently, we propose a novel reward design mechanism based on robot CoT and sparse failure guidance.
-Specifically, we explore the usage of VideoCLIP (VideoCLIP) as a vision-language foundation model.
+In this work, we argue that VLMs can provide the agent with a much more fine-grained reward guidance.
+Consequently, we propose a novel reward design mechanism based on robot CoT.
+Specifically, we explore the usage of VideoCLIP [29] as a VLM.
 VideoCLIP excels in capturing the temporal dynamics of video segments and interpreting associated language prompts, aiming to align robot movements with actions described in the task descriptions.
@@
 $\text{reward}=\frac{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)}{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (3)
-where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task descriptions from robot CoT interpretation, and $z_{l}^{-}$ is a set of potential failure case descriptions guided by humans.
+where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task description embeddings from robot CoT interpretation $l^{+}$, and $z_{l}^{-}$ is a set of potential failure case description embeddings guided by humans.
+Here, offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
+To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training with 2 to 4 prompts.
 The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims to distinguish between video segments and corresponding positive or negative texts.
-We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP (RoboCLIP).
-Moreover, the max operation in Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent.
+We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP [2].
+Moreover, Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent.
 Otherwise, the reward signal would penalize undesired behaviors.
-4.3 Policy Learning with Success Experience Regularization
+4.3 VLM-based Self Imitation
 Our reward design mechanism motivates the agent to explore the necessary manipulation skills.
-Although this reward signal performs well in manipulation tasks, the VideoCLIP rewards do not reflect the completion of the manipulation tasks.
-They only indicate whether the robot’s movements align with detailed language goals.
-To better reinforce task completion, we introduce policy learning with success experience regularization.
+Empirically, we find that the VideoCLIP rewards don’t reflect the completion of the manipulation tasks well.
+They focus mostly on whether the robot’s movements align with detailed language goals.
+To better reinforce task completion and speed up the learning, we introduce VLM-based self imitation procedure.
 We employ GPT-4V to assess the success of the agent’s final observation in its rollout trajectory.
-If it deemed successful, this trajectory is archived into a success experience replay buffer, enabling the agent to reinforce this success.
+If it deemed successful, this trajectory is archived into a success replay buffer and we correct the reward function as: $r(s_{T})=r(s_{T})+100$.
+Here, $s_{T}$ is the successful final observation of the rollout trajectory.
 Subsequently, we introduce a regularization mechanism aimed at refining the agent’s policy to reinforce successful behavior.
+This motivates the agent to reinforce the success experience.
 The regularization loss is integrated with the policy loss, formulated as follows:
@@
 By incorporating success experiences into the policy loss calculation, the agent is encouraged to repeat desired behaviors.
-However, the reward signal for success experiences should also be reinforced to help the agent understand the task completion signal.
-Therefore, we correct the reward function as follows: $r(s_{T})=r(s_{T})+100$, where $s_{T}$ is the successful final observation of the rollout trajectory.
-4.4 Pipeline
-Our algorithm begins with GPT-4V interpreting language instructions into detailed task descriptions.
-Meanwhile, humans guide the agent with several potential failure case descriptions.
-Subsequently, rewards are calculated following Equation 3 with a moving window that slides over the agent’s rollout trajectories.
-The agent is then updated with these learned rewards to further develop the skills.
-Concurrently, GPT-4V evaluates the final observation of the rollout trajectory and assigns a preference label, indicating whether it is a successful observation.
-If it is successful, we record the trajectory into a success experience replay buffer for policy regularization and correct the reward function.
-Figure 2 illustrates our method for computing rewards and success regularization that aid the robot in learning manipulation skills.
-This involves a moving window that traverses the agent’s rollout trajectory, the computation of NCE rewards using language prompts, and the policy regularization using success experiences gathered through GPT-4V.
-This comprehensive approach aims to effectively teach the agent desired manipulation skills through reinforcement learning.
 5 Experiment
+In this section, we assess the effectiveness of our proposed algorithm RoboCoT.
+The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms?
+The results are presented in section 5.1.
+(2) What are the contributing factors in our algorithm?
+The ablation studies are presented in section 5.2.
+Environments To evaluate the effectiveness of our proposed algorithm RoboCoT, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment.
+We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by [30].
+Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
+Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 [31] as the reinforcement learning algorithm.
+Detailed implementation specifics can be found in the references: CLIP [24], LIV [25], and RoboCLIP [2].
+All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs.
+Implementation details are provided in the Appendix 8.
+5.1 Overall Comparison
 Figure 3: Training curves for baselines and our algorithm within 3 random seeds.
@@
 Our algorithm gets best performance compared to other baselines, achieves 5.4×5.4$\times$5.4 × improvement compared to RoboCLIP.
-In this section, we assess the effectiveness of our proposed algorithm by conducting experiments in comparison with several baseline models.
-The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms?
-The results are presented in section 5.1.
-(2) What are the contributing factors in our algorithm?
-The ablation studies are presented in section 5.2.
-Environments To evaluate the effectiveness of our proposed algorithm, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment.
-We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by (ROT).
-Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
-Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 (DrQv2) as the reinforcement learning algorithm.
-Detailed implementation specifics can be found in the references: CLIP (MetaCLIP), LIV (LIV), and RoboCLIP (RoboCLIP).
-All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs.
-Implementation details are provided in the Appendix.
-5.1 Overall Comparison
-To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP (MetaCLIP), LIV (LIV), and RoboCLIP (RoboCLIP).
-CLIP (as described in (MetaCLIP)) employs CLIP (CLIP) as the vision-language model to measure the distance between current observations and language goals.
-LIV (LIV) aligns goal representations with language instructions, and RoboCLIP (RoboCLIP) utilizes S3D (S3D) to measure the distance between trajectory videos and language as a form of reward.
-However, CLIP (CLIP) and LIV (LIV) lack an understanding of the manipulation process, while RoboCLIP (RoboCLIP) lacks spatial reasoning capabilities and task completion awareness.
+To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP [24], LIV [25], and RoboCLIP [2].
+CLIP [24] employs CLIP [23] as the vision-language model to measure the distance between current observations and language goals.
+LIV [25] aligns goal representations with language instructions, and RoboCLIP [2] utilizes S3D [3] to measure the distance between trajectory videos and language as a form of reward.
+However, CLIP [23] and LIV [25] lack an fine-grained reward guidance, while RoboCLIP [2] lacks task completion awareness.
 Figure 3 illustrates the training curves, depicting the success rates for each algorithm.
@@
 LIV shows competence primarily in tasks involving drawer closure, window opening, and soccer.
-Our algorithm consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors.
+Our algorithm RoboCoT consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors.
 These results underscore the effectiveness of our approach in handling diverse and challenging scenarios.
 5.2 Ablation Study
-In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and success regularization).
-The results indicate that the most contributing factor to our algorithm’s performance is success regularization, followed by human failure guidance.
+In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and self imitation).
+The results indicate that the most contributing factor to our algorithm’s performance is VLM-based self imitation procedure, followed by human failure guidance.
 Each component contributes to the overall performance of our algorithm.
 Additionally, we conduct ablation studies comparing GPT-4V success labels with ground truth success labels, demonstrating the robustness of our GPT-4V labeling.
-Furthermore, we explore different reward design mechanisms, showing that they can also achieve high performance within our algorithm.
-Finally, experiments with all baselines incorporating success regularization reveal that our algorithm’s superiority is not solely due to success regularization but also the reward design mechanism.
+Finally, experiments with all baselines incorporating VLM-based self imitation procedure reveal that our algorithm’s superiority is not solely due to self imitation but also the reward design mechanism.
 Question 1: how significant is each component of our algorithm in contributing to the results?
-Figure 4: Final policy performance for ablations and our algorithm across 3 random seeds.
-The results show that success regularization is the most significant contributing factor in our algorithm, followed by human failure guidance.
-Each component contributes to the overall performance of our algorithm.
-Our algorithm incorporates several critical components: robot chain-of-thought processing via GPT-4V, sparse human failure guidance, and success experience regularization.
+Figure 4: Final policy performance for ablations and our algorithm.
+The results show that VLM-based self imitation procedure is the most significant contributing factor in our algorithm, followed by human failure guidance.
+Our algorithm incorporates several critical components: robot chain-of-thought, sparse human failure guidance, and VLM-based self imitation.
 To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance.
@@
 Figure 4 illustrates the success rates of the final policy performance of our algorithm compared to its ablated versions.
-The labels for these ablations are: w/oReg (without success experience regularization), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
-The results confirm the critical importance of each component in our algorithm, with success experience regularization being particularly pivotal.
+The labels for these ablations are: w/oReg (without success experience regularization, i.e., self imitation), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
+The results confirm the critical importance of each component in our algorithm, with VLM-based self imitation being particularly pivotal.
 Human failure guidance is the second most important component, motivating the agent to consider potential failure cases.
@@
 The results demonstrate that the success labels generated by GPT-4V are robust.
-Figure 6: Ablation study for our algorithm with CoT as one single prompt or multiple prompts. The results show that single prompt for reward mechanism also works well.
-Our algorithm incorporates GPT-4V to gather successful trajectories for policy regularization.
+Figure 6: Ablation study for all baselines with success experience regularization.
+The results show that our algorithm performs well not only because of the VLM-based self imitation component but also due to the reward design mechanism.
+Our algorithm RoboCoT incorporates GPT-4V to gather successful trajectories for VLM-based self imitation procedure.
 To validate the robustness of our GPT-4V labeling in collecting success experiences, we conduct experiments comparing GPT-4V labels and ground truth labels.
@@
 Otherwise, the output is 0." Subsequently, we present the final observation to GPT-4V and prompt, "What is the output of this image?" We collect the trajectory as successful when GPT-4V’s response is 1.
-Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labeling methods for success experience collection.
-Our findings show that the GPT-4V labeling method achieves performance comparable to the ground truth across various tasks, demonstrating the robustness of our GPT-4V success labels.
-Question 3: what is the empirical difference between our algorithm and the one with a single chain-of-thought interpreted positive prompt?
-Our algorithm computes the NCE-rewards based on the maximum distance between the moving window video and multiple positive prompts.
-This measurement indicates how far the moving window video is from the language goal derived from robot chain-of-thought interpretations.
-Conversely, our ablation method computes the NCE-rewards based on the distance between the moving window video and a single interpreted positive prompt.
-The NCE-reward function is shown as:
-$\text{reward}=\frac{\exp(z_{v}\cdot z_{l}^{+}/\tau)}{\exp(z_{v}\cdot z_{l}^{+}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (5)
-where $z_{l}^{+}$ indicates the single positive prompt from the GPT-4V interpretation.
-For example, the prompt for opening a door is "the robot hand first approaches the door’s handle, then grasps it, pulls it back, and finally opens the door."
-Figure 6 shows the comparison between our algorithm with multiple positive prompts and the single positive prompt ablation.
-The empirical results demonstrate that our algorithm achieves comparable results with a single positive prompt.
-It indicates that our algorithm is robust to the setting of multiple or single positive prompts.
-Figure 7: Ablation study for all baselines with success experience regularization.
-The results show that our algorithm performs well not only because of the success regularization component but also due to the reward design mechanism.
-Question 4: what is the empirical difference between our algorithm and all baselines with success experience regularization?
-Question 1 shows that success experience regularization component plays Question 1 shows that the success experience regularization component plays the most important role in our algorithm.
-To evaluate the significance of the reward design mechanism, we conduct experiments applying success regularization to all baselines.
-If the results show that our algorithm still outperforms other baselines with success regularization, then it proves that our algorithm’s superiority is due not only to the success regularization but also to the reward design mechanism.
-Figure 7 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include success experience regularization.
+Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labels for success experience collection.
+Our findings show that the GPT-4V labels achieves comparable performance to the ground truth across various tasks, demonstrating the robustness of our GPT-4V labels.
+Question 3: what is the empirical difference between our algorithm and all baselines with VLM-based self imitation?
+Question 1 shows that VLM-based self imitation component plays the most important role in our algorithm.
+To evaluate the significance of the reward design mechanism, we conduct experiments applying VLM-based self imitation procedure to all baselines.
+If the results show that our algorithm still outperforms other baselines with self imitation procedure, then it proves that our algorithm’s superiority is due not only to the VLM-based self imitation procedure but also to the reward design mechanism.
+Figure 6 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include VLM-based self imitation procedure.
 The results demonstrate that our algorithm outperforms other baselines on almost all tasks.
-This indicates that our algorithm’s superiority is due not only to the success regularization component but also to the reward design mechanism.
+This indicates that our algorithm’s superiority is due not only to the VLM-based self imitation component but also to the reward design mechanism.
 6 Conclusion
-In this study, we present a novel algorithm for robotic manipulation skill acquisition.
-We propose that pre-trained vision-language models can instruct robots on prior knowledge essential for manipulation tasks.
-Our algorithm simplifies complex tasks and acquires manipulation skills through a robot’s chain-of-thought process with sparse failure guidance, mirroring the developmental learning trajectory of a child.
-Specifically, our algorithm is designed to extract intricate temporal visual cues from the robot’s thought process, enabling a nuanced understanding of the task at hand.
-It also incorporates a learning strategy by describing potential failure cases through sparse failure guidance, which enhances its manipulation policies.
-Through success experience regularization, our algorithm strengthens its policies, thereby improving performance.
+In this study, we present a novel algorithm RoboCoT, which defines rewards for skill acquisition.
+Our algorithm RoboCoT simplifies complex tasks through a robot’s chain-of-thought process with sparse failure guidance.
+It yields a more fine-grained reward guidance.
+Through VLM-based self imitation, our algorithm strengthens its policies, thereby improving performance.
 Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage over established baseline models.
-This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic manipulation tasks in the future.
-7 Appendix
-7.1 Environment Setup
-We use MetaWorld-v2 as our test robotic environments.
-We choose 10 tasks as by [ROT] to evaluate our algorithm and baselines.
-They are drawer close, window close, door close, faucet close, button press, drawer open, window open, door open, faucet open and soccer.
-Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
-To reduce the exploration space, we set the gripper a new position.
-With varying random seeds, the objects like drawer, window, door, faucet, button, and soccer would vary its positions.
-The position coordinate are provided in Table 1.
-Table 1: Position coordinates for robot gripper as setting with MetaWorld-v2.
-Task Coordinate
-Drawer Close $0,0.6,0.2$
-Drawer Open $0,0.8,0.05$
-Window Close $0.25,0.80,0.16$
-Window Open $-0.15,0.77,0.16$
-Door Close $-0.5,0.6,0.2$
-Door Open $0,0.6,0.2$
-Faucet Close 0.,0.6,0.10.,0.6,0.10 . , 0.6 , 0.1
-Faucet Open 0.,0.6,0.10.,0.6,0.10 . , 0.6 , 0.1
-Button Press $0,0.6,0.2$
-Soccer $0.6,0$ 0.6 , 0.06
-Since the default camera view is not suitable for our algorithms.
-We re-set the camera for training all the experiments.
-The principle for choosing camera is that we set the object and gripper in the view.
-The camera coordinates are provided in Table 2.
-Table 2: Position coordinates and angles for camera as setting with MetaWorld-v2.
-Task Coordinate Angle
-Door Close $-1.1,-0.4,0.6$
-Other Tasks $0.55,0.3,0.5$ $3.9,2.3,0.6$
-7.2 GPT-4V Success Labeling Details
-Figure 8: The successful trajectories are gathered through GPT-4V success labeling.
-GPT-4V would label the success preference to the final observation of agent’s rollout trajectory.
-If it deemed successful, the trajectory would be recorded into a replay buffer.
-Our algorithm incorporates GPT-4V to gather successful trajectories for policy regularization.
-The process for GPT-4V analysis is outlined in Figure 8.
-Initially, we present GPT-4V with a scenario indicating a successful state (e.g., the door is open).
-We then provide a prompt: "Consider the following image; this is a simulation environment.
-In this image, the door of the box is open.
-We define it when the door is open, the output is 1.
-Otherwise, the output is 0." Subsequently, we present the final observation to GPT-4V and prompt, "What is the output of this image?" We collect the trajectory as successful when GPT-4V’s response is 1.
-To reduce the query times with GPT-4V, we distill a classifier from GPT-4V labeling data.
-First, we use GPT-4V to label 300 pictures with diverse final observations.
-Next, we fine-tune a ResNet-18 classifier based on this dataset.
-Before training, we use some augmentation techniques such as random crop resize to augment the dataset.
-Finally, we use this classifier to classify the final observation as successful or not for instance.
-7.3 Task Robot Chain-of-Thought Prompt
-Manipulation skills consist of a series of actionable control steps.
-For instance, opening a door includes sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door.
-Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control.
-We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $z_{l}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
-The task robot CoT prompts are provided in Table 3.
-Table 3: Prompts from robot chain-of-thought for task interpretation.
-Task Robot Chain-of-Thought Prompts
-Drawer Close the robot hand pushes the drawer;
-the drawer is closed.
-Drawer Open the robot hand moves to the top of the drawer’s handle;
-the robot hand moves inside the drawer’s handle;
-the robot hand pulls back the drawer’s handle;
-the robot hand opens the drawer.
-Window Close the robot hand moves to the left of the window’s handle;
-the robot hand pushes the handle towards right;
-the robot hand closes the window.
-Window Open the robot hand moves to the right of the window’s handle;
-the robot hand pushes the handle towards left;
-the robot hand opens the window.
-Door Close the robot hand pushes the door close;
-the door is closed.
-Door Open the robot hand is at the front of the door;
-the robot hand moves to the left of the door’s handle;
-the robot hand pulls back the door’s handle;
-the robot hand opens the door
-Faucet Close the robot hand moves to the left of the faucet;
-the robot hand pulls the faucet towards right;
-the faucet is closed.
-Faucet Open the robot hand moves to the right of the faucet;
-the robot hand pulls the faucet towards left;
-the faucet is open.
-Button Press the robot hand moves towards the red button;
-the robot hand pushes the button;
-the button is pressed.
-Soccer the robot hand moves towards soccer;
-the robot hand pushes soccer into the net;
-the soccer is in the net.
-7.4 Sparse Human Failure Guidance
-Robots may encounter potential failures during learning.
-Offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
-$z_{l}^{-}$ is a set of offline potential failure case descriptions consisting of 2 to 4 prompts.
-To enable the agent to recognize and adjust to potential failures, we provide sparse human failure guidance before training the agent.
-The sparse human failure guidance are provided in Table 4.
-Table 4: Sparse human failure guidance.
-Task Human Failure Guidance
-Drawer Close the robot hand is moving away from the drawer;
-the robot hand is at the left of the drawer;
-the robot hand is at the right of the drawer.
-Drawer Open the robot hand is at the right of the drawer’s handle;
-the robot hand is at the left of the drawer’s handle;
-the robot hand is under the drawer’s handle.
-Window Close the robot hand is at the top of the window’s handle;
-the robot hand is at the right of the window’s handle;
-the robot hand is under the window’s handle.
-Window Open the robot hand is at the top of the window’s handle;
-the robot hand is at the left of the window’s handle;
-the robot hand is under the window’s handle.
-Door Close the robot hand is moving away from the door;
-the robot hand is at the left of the door.
-Door Open the robot hand is at the right of the door’s handle;
-the robot hand is at the top of the door’s handle;
-the robot hand is under the door’s handle;
-the robot hand is at the left of the box.
-Faucet Close the robot hand is to the right of the faucet;
-the robot hand is at the top of the faucet.
-Faucet Open the robot hand is to the left of the faucet;
-the robot hand is at the top of the faucet.
-Button Press the robot hand moves away from the red button;
-the robot hand is at the left of the red button;
-the robot hand is at the right of the red button.
-Soccer the robot hand is moving away from soccer;
-the robot hand is at the top of soccer.
+This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic tasks in the future.
+7 Limitation
+Our proposed method, RoboCoT, effectively masters a wide range of manipulation tasks through a robot chain-of-thought with sparse failure guidance.
+It provides a fine-grained reward guidance for the robot agent, facilitating the training of various automation tools.
+However, these tasks primarily focus on aligning robot movements with detailed language descriptions.
+RoboCoT does not yield promising results in tasks that emphasize observational preferences, such as soccer.

### v3

@@
 Recently, vision-language models (VLMs) have shown promise in defining reward signals for teaching robots manipulation skills.
-However, existing works often provide reward guidance that is too coarse, leading to inefficient learning processes.
+However, existing work often provides reward guidance that is too coarse, leading to insufficient learning processes.
 In this paper, we address this issue by implementing more fine-grained reward guidance.
@@
 Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP.
-Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rate compared to the best baseline, RoboCLIP, across a series of manipulation tasks.
-Keywords: Manipulation Skills, Vision Language Models, Chain-of-Thought
-1 Introduction
-Defining reward functions for learning robotic skills has been a long standing challenge in robotics [1].
-Traditional methods typically define the rewards through human knowledge and require significant engineering efforts.
+Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rates compared to the best baseline, RoboCLIP, across a series of manipulation tasks.
+I INTRODUCTION
+Defining reward functions for learning robotic skills has been a long-standing challenge in robotics [1].
+Traditional methods typically define rewards through human knowledge and require significant engineering efforts.
 Some methods also use expert demonstrations to define rewards [1].
 However, this procedure requires expert supervision and can be infeasible in complex scenarios [2] .
-Recently, the success of vision language models (VLM) has inspired researchers to leverage its power to define rewards for skill learning.
-For example, RoboCLIP [2] captures the temporal information from robot movements with a video backbone S3D [3].
+Recently, the success of vision language models (VLM) has inspired researchers to use its power to define rewards for skill learning.
+For example, RoboCLIP [2] captures the temporal information from robot movements with an S3D video backbone [3], therefore defining the rewards for skill learning.
 In spite of the preliminary success of VLM-based reward modeling, a downside of these existing works is that their language guidance is too coarse for learning.
 Consider the task of opening the door, existing works only use the language embedding of "opening the door" in reward definition to guide the learning process.
-Such guidance can be too sparse for learning: when the robot is not able to approach the door at the beginning of learning process, the rewarder will always provide a near-zero feedback since the door is not likely to be opened.
+Such guidance can be too coarse for learning: When the robot is not able to approach the door at the beginning of learning process, the rewarder will always provide a near-zero feedback since the door is not likely to be opened.
 In contrast, in human world we typically use more fine-grained guidance such as "move towards the door", "move the hand towards the handle" and "grab the handle" during the teaching process.
-Intuitively, using such decomposition can make the guidance more precise, and will lead to richer and denser feedback.
-Based on this intuition, in this paper we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RoboCoT) to address this problem.
+Intuitively, using such decomposition [4, 5, 6, 7, 8] can make the guidance more precise, and will lead to richer and denser feedback.
+Based on this intuition, in this paper, we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RoboCoT) to address this problem.
 Our main insight is to decompose the task description into a series of actionable steps and then leverage VLM to define a fine-grained reward to guide skill learning, which can make the learning process more sample efficient and yield higher success rates.
-We also propose a VLM-based self-imitation procedure to further speed up the learning process.
+In addition, we propose a VLM-based self-imitation procedure to further speed up the learning process.
 In the experiments, we demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline.
-Ablation studies show that each component of our algorithm contributes to the overall performance.
+Empirical studies show that each component of our algorithm contributes to the overall performance.
 Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using a robot chain-of-thought processing, breaking it down into several detailed prompts.
-The robot also considers some failure guidance.
-Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills.
-Successful experiences are recorded by foundation models to reinforce effective behaviors (i.e., VLM-based self-imitation).
-2 Related Work
-2.1 Learning Manipulation Skills for Robotics
-Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) [4, 5, 6] or inverse reinforcement learning (IRL) [7, 8].
-BC directly maps the state space to the action space to mimic the expert’s actions.
-However, BC faces several limitations, including compounding errors [9], the copycat problem [10], and out-of-distribution sample shifts [11].
-In contrast, IRL derives both the reward function and the policy from expert demonstrations.
-However, IRL struggles with sample efficiency, particularly in real-world applications [12, 13, 14].
-Mastering manipulation skills with language inputs directly presents several challenges, such as a lack of reward signals and inefficiency in understanding vision-language contexts.
-Recent studies have explored integrating VLMs to enhance the understanding of language instructions [15].
-Our algorithm is designed by effectively leveraging VLMs as a source of rewards.
-2.2 Foundation Models for Robotics
-VLMs equip robot agents with compact pre-trained representations.
+Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills with the consideration about the failure guidance.
+Successful experiences are recorded by foundation models to reinforce effective behaviors, i.e., VLM-based self-imitation.
+II RELATED WORK
+Defining reward functions for robotic skill learning has long been a significant challenge.
+This challenge can be addressed through learning from demonstrations using inverse reinforcement learning (IRL) [9, 10, 11, 12, 13].
+However, IRL requires expert supervision, which can be infeasible in complex scenarios [2].
+Our RoboCoT overcomes this limitation by effectively leveraging VLMs as a source of rewards.
+By fully utilizing prior knowledge from VLMs, RoboCoT comprehends the manipulation process and subsequently achieves superior manipulation performance by learning directly from language instructions.
+II-A Robot Chain-of-Thought for Robotics
+Robot chain-of-thought [4] is employed to interpret task descriptions and enhance high-level planning performance in robotics.
+For example, ViLa [7] utilizes GPT-4V to decompose tasks into sub-goals based on robot observations and task descriptions, demonstrating significant success in understanding complex robotic tasks.
+HiP [5] achieves hierarchical planning by diffusing the future observations, therefore better decomposing the complex tasks.
+However, current approaches primarily apply robot chain-of-thought to high-level planning, thereby limiting its potential for comprehensively understanding robotic tasks.
+Our RoboCoT extends the application of robot chain-of-thought by designing more fine-grained reward signals.
+Empirical studies demonstrate its promise in directly following language instructions.
+II-B Vision Language Models for Robotics
+VLMs equip robot agents with compact, pre-trained representations.
 R3M [14] co-trains video and language instructions in a self-supervised manner [15], enhancing the robot’s ability to mimic expert actions.
@@
 RT-2 [19] innovates by conceptualizing robot actions as a form of language.
-Furthermore, RT-H [22] introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large language models.
+RT-H [20] introduces an action hierarchy through language motions, leveraging the robot chain-of-thought reasoning capabilities of large language models (LLMs).
 VLMs significantly enhance robot manipulation capabilities by offering valuable reward signals.
@@
 RoboCLIP [2] adopts the S3D [3] representation model to assess the reward for the final observation of a trajectory.
-Furthermore, RL-VLM-F [27] directly utilizes GPT-4V to label preferences for agent rollout data.
-However, current models still struggle to master a broad spectrum of robot tasks, often succeeding only in simple tasks such as closing a drawer.
-VLMs like CLIP [23] and LIV [25] lack an understanding of the manipulation process.
-On the other hand, RoboCLIP [2] lacks spatial reasoning capabilities and task completion awareness.
-RL-VLM-F [27] excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities.
-In contrast, our algorithm not only comprehends detailed language instructions from the robot chain-of-thought but also considers failure scenarios.
-Additionally, it leverages VLM’s task completion reasoning ability through self imitation process, offering a more comprehensive approach to master robot tasks.
-3 Background
+Additionally, [25] considers potential failure cases for designing reward signals in imitating language instructions.
+RL-VLM-F [26] directly uses GPT-4V to label preferences for agent rollout data.
+However, current models use coarse instructions like "open the door" for designing reward signals.
+Our RoboCoT improves upon this by leveraging the robot chain-of-thought technique to design more fine-grained reward signals for following language instructions.
+Empirical studies demonstrate the promise of using VLMs as rewarders in imitation learning.
+III BACKGROUND
 We consider scenarios where the robotic system receives specific verbal commands, such as "open the door." To model this learning process, we employ a Markov Decision Process (MDP) defined by $(S,A,r,P,\gamma)$, where $S$ denotes the state space, $A$ the action space, $r$ the reward model, $P$ the transition model, and $\gamma$ the discount factor.
-Chain-of-Thought (CoT) The CoT [28] approach is a reasoning method designed to enhance the problem-solving capabilities of large language models.
+Chain-of-Thought (CoT) The CoT [4] approach is a reasoning method designed to enhance the problem-solving capabilities of LLMs.
 This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving.
-By explicitly detailing these intermediate steps, large language models such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
+By explicitly detailing these intermediate steps, LLMs such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
 VideoCLIP VideoCLIP [27] is a multi-modal model that connects visual and textual information, specifically designed to understand videos by associating them with relevant text descriptions.
@@
 where $\tau$ is a temperature hyper-parameter, $z_{t}^{+}$ and $z_{t}^{-}$ are positive and negative text embeddings respectively, and $z_{v}$ is a video segment embedding.
-4 Method
-Figure 2: First, text embeddings are generated using robot chain-of-thought (positive prompts) and human failure guidance (negative prompts).
+IV METHOD
+Figure 2: First, text embeddings are generated with robot chain-of-thought (positive prompts) and sparse failure guidance (negative prompts).
 Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings.
 Finally, VLMs evaluate the final observation.
-If this observation is deemed successful, this trajectory is recorded in the success replay buffer for self imitation.
-In section 4.1, we introduce the robot chain-of-thought interpretation.
-In section 4.2, we explain how to define the rewards for policy learning with VLMs.
-Finally, in section 4.3, we introduce VLM-based self imitation procedure via success experience regularization.
-4.1 Robot Chain-of-Thought
-Manipulation skills consist of a series of actionable control steps.
-For instance, opening a door includes sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door.
-Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control.
-We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $l_{i}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
-4.2 VLMs as Rewarder
-A central challenge we address is defining a reward function that effectively reinforces desired behaviors.
-In this work, we argue that VLMs can provide the agent with a much more fine-grained reward guidance.
-Consequently, we propose a novel reward design mechanism based on robot CoT.
-Specifically, we explore the usage of VideoCLIP [29] as a VLM.
-VideoCLIP excels in capturing the temporal dynamics of video segments and interpreting associated language prompts, aiming to align robot movements with actions described in the task descriptions.
-The reward signal for learning manipulation skills is defined as follows:
-$\text{reward}=\frac{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)}{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (3)
-where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task description embeddings from robot CoT interpretation $l^{+}$, and $z_{l}^{-}$ is a set of potential failure case description embeddings guided by humans.
-Here, offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills.
-To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training with 2 to 4 prompts.
-The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims to distinguish between video segments and corresponding positive or negative texts.
-We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP [2].
+If this observation is deemed successful, this trajectory is recorded in the replay buffer for self imitation procedure.
+In Section IV-A, we present the robot chain-of-thought process for comprehending the robotic tasks.
+In Section IV-B, we outline the process of defining rewards for skill learning using Vision-Language Models (VLMs).
+Finally, in Section IV-C, we introduce the VLM-based self-imitation procedure through success experience regularization.
+IV-A Robot Chain-of-Thought
+Robotic manipulation skills consist of a series of actionable control steps.
+For example, opening a door involves sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door.
+Decomposing tasks with a robot chain-of-thought (CoT) process simplifies the manipulation process, providing robots with a deeper understanding and greater control.
+However, previous works [2, 25, 22, 23] have primarily used coarse language instructions, such as ’open the door,’ in defining reward signals.
+This level of guidance is insufficient for learning robust manipulation skills.
+Our RoboCoT addresses this limitation by employing robot chain-of-thought process to decompose tasks for more fine-grained reward design.
+We use GPT-4V to interpret task instructions like ’open the door.’ Through robot chain-of-thought process, GPT-4V breaks down this instruction into a series of sub-goals, such as ’the robot hand approaches the door’s handle,’ ’the robot hand grasps the door’s handle,’ ’the robot hand pulls the door’s handle back,’ and ’the robot hand opens the door.’ Consequently, we leverage these task descriptions from the GPT-4V interpretation to define a fine-grained reward signal in skill learning, thereby enhancing the agent’s understanding of the robotic tasks.
+IV-B VLMs as Rewarder
+Defining reward functions for robot skill learning is a long standing challenge.
+In this work, we leverage VLMs to design more fine-grained reward signals in skill learning.
+Moreover, we also consider the potential failure guidance in teaching robot manipulation skills which has been investigated in [25].
+We explore the usage of VideoCLIP [27] as a rewarder.
+VideoCLIP excels in capturing the temporal dynamics of video slices and interpreting associated language prompts, aiming at aligning robot movements with actions described in the task descriptions.
+We define the reward signal for skill learning as:
+$r=\frac{\max\limits_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)}{\max\limits_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)+\sum\limits_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (3)
+where $r$ represents the rewards for moving window stride timesteps, $z_{v}$ is the video embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task description embeddings from robot chain-of-thought process, and $z_{l}^{-}$ is a set of potential failure case embeddings guided by humans.
+Here, sparse failure guidance motivates the agent to consider potential failure cases, improving the robustness.
+The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims at distinguishing between video slices and corresponding positive or negative texts.
+We find that this similar NCE-based reward formulation can better leverage the VideoCLIP model which is different from the reward definition in RoboCLIP [2].
 Moreover, Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent.
-Otherwise, the reward signal would penalize undesired behaviors.
-4.3 VLM-based Self Imitation
-Our reward design mechanism motivates the agent to explore the necessary manipulation skills.
-Empirically, we find that the VideoCLIP rewards don’t reflect the completion of the manipulation tasks well.
-They focus mostly on whether the robot’s movements align with detailed language goals.
-To better reinforce task completion and speed up the learning, we introduce VLM-based self imitation procedure.
-We employ GPT-4V to assess the success of the agent’s final observation in its rollout trajectory.
-If it deemed successful, this trajectory is archived into a success replay buffer and we correct the reward function as: $r(s_{T})=r(s_{T})+100$.
-Here, $s_{T}$ is the successful final observation of the rollout trajectory.
-Subsequently, we introduce a regularization mechanism aimed at refining the agent’s policy to reinforce successful behavior.
-This motivates the agent to reinforce the success experience.
-The regularization loss is integrated with the policy loss, formulated as follows:
+Figure 3: Training curves for baselines and our algorithm within 3 random seeds.
+Each data point is evaluated with 20 sampled trajectories.
+The shaded area displays the range of one standard deviation.
+Our algorithm gets best performance compared to other baselines, achieves 5.4×5.4$\times$5.4 × average improvement compared to RoboCLIP.
+IV-C VLM-based Self Imitation
+Our reward design mechanism encourages the agent to explore the necessary skills.
+We find that the VideoCLIP rewards focus on whether the robot movements align with the detailed language instructions.
+Consequently, speeding up the learning is necessary.
+To better speed up the learning, we introduce VLM-based self imitation procedure.
+We employ VLMs, i.e., GPT-4V, to assess the success of the agent’s final observation in its rollout trajectory.
+If it deemed successful, this trajectory is recorded into a success replay buffer $D_{E}$.
+Additionally, we correct the reward function on this successful observation as: $r(o_{T})=r(o_{T})+100.0$.
+Here, $r(o_{T})$ is the reward function for this successful final observation $o_{T}$ in the rollout trajectory.
+Subsequently, we introduce a regularization term in loss function, aimed at reinforcing success behaviors.
+The regularization loss is integrated with the policy RL loss, formulated as follows:
 =RL $\mathcal{L}=\text{RL Loss}+\lambda\sum_{(s,a)\in D_{E}}-\log(\pi(a|s)),$ (4)
 where $D_{E}$ is the success experience replay buffer and $\lambda$ is the hyper-parameter.
-By incorporating success experiences into the policy loss calculation, the agent is encouraged to repeat desired behaviors.
-5 Experiment
+By incorporating success experiences into the policy loss calculation, we speed up the skill learning process.
+Here we note that in our experiments, $\lambda$ is set to $1$.
+V EXPERIMENT
 In this section, we assess the effectiveness of our proposed algorithm RoboCoT.
-The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms?
-The results are presented in section 5.1.
-(2) What are the contributing factors in our algorithm?
-The ablation studies are presented in section 5.2.
+The experiments are designed to address two questions: (1) How does our algorithm RoboCoT perform compared to other state-of-the-art algorithms?
+The results are presented in section V-A.
+(2) What are the contributing factors in our RoboCoT?
+The ablation studies are presented in section V-B.
 Environments To evaluate the effectiveness of our proposed algorithm RoboCoT, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment.
-We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by [30].
+We select 10 tasks from simple to complex, including drawer, window, faucet, button, door, and soccer, as described in [28].
 Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
@@
 Detailed implementation specifics can be found in the references: CLIP [22], LIV [23], and RoboCLIP [2].
-All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs.
-Implementation details are provided in the Appendix 8.
-5.1 Overall Comparison
-Figure 3: Training curves for baselines and our algorithm within 3 random seeds.
-Each data point is evaluated with 20 sampled trajectories.
-The shaded area displays the range of one standard deviation.
-Our algorithm gets best performance compared to other baselines, achieves 5.4×5.4$\times$5.4 × improvement compared to RoboCLIP.
-To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP [24], LIV [25], and RoboCLIP [2].
-CLIP [24] employs CLIP [23] as the vision-language model to measure the distance between current observations and language goals.
-LIV [25] aligns goal representations with language instructions, and RoboCLIP [2] utilizes S3D [3] to measure the distance between trajectory videos and language as a form of reward.
-However, CLIP [23] and LIV [25] lack an fine-grained reward guidance, while RoboCLIP [2] lacks task completion awareness.
+All the experiments, including those involving our proposed algorithm RoboCoT, receive the same language instructions as inputs.
+V-A Overall Comparison
+To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing our RoboCoT with existing baselines: CLIP [22], LIV [23], and RoboCLIP [2].
+CLIP [22] employs CLIP [21] as the VLMs to measure the distance between current observations and language goals.
+LIV [23] aligns VIP [24] goal representations with language instructions, therefore measuring the reward distance between current robot observation with language goals.
+RoboCLIP [2] utilizes S3D [3] to measure the distance between trajectory videos and language as a form of reward.
+Since the reward is defined with a coarse language instruction, CLIP [21], LIV [23] and RoboCLIP [2] lack fine-grained reward guidance.
+On the other hand, our RoboCoT uses robot chain-of-thought and sparse failure guidance to define the rewards for skill learning.
+Consequently, RoboCoT can better acquire the skills through reinforcement learning.
 Figure 3 illustrates the training curves, depicting the success rates for each algorithm.
 Each evaluation involves 20 trajectories, and the experiments were replicated across three different environmental random seeds to introduce variability in the position of objects.
+Figure 4: Final policy performance for ablations and our RoboCoT. The results show that each component of our RoboCoT contributes to the final performance.
 Our findings indicate distinct performance trends across various tasks.
-RoboCLIP displays promising results in tasks involving drawer and window manipulation, as well as soccer, but it is most effective in the simplest tasks such as drawer closing.
-CLIP performs well in tasks related to faucet, drawer and window manipulation, and soccer, indicating that CLIP focuses more on task completion rather than the manipulation process.
-LIV shows competence primarily in tasks involving drawer closure, window opening, and soccer.
-Our algorithm RoboCoT consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors.
-These results underscore the effectiveness of our approach in handling diverse and challenging scenarios.
-5.2 Ablation Study
-In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and self imitation).
-The results indicate that the most contributing factor to our algorithm’s performance is VLM-based self imitation procedure, followed by human failure guidance.
-Each component contributes to the overall performance of our algorithm.
-Additionally, we conduct ablation studies comparing GPT-4V success labels with ground truth success labels, demonstrating the robustness of our GPT-4V labeling.
-Finally, experiments with all baselines incorporating VLM-based self imitation procedure reveal that our algorithm’s superiority is not solely due to self imitation but also the reward design mechanism.
-Question 1: how significant is each component of our algorithm in contributing to the results?
-Figure 4: Final policy performance for ablations and our algorithm.
-The results show that VLM-based self imitation procedure is the most significant contributing factor in our algorithm, followed by human failure guidance.
-Our algorithm incorporates several critical components: robot chain-of-thought, sparse human failure guidance, and VLM-based self imitation.
+RoboCLIP displays promising results in the simplest tasks.
+CLIP performs well in tasks related to faucet, drawer and window manipulation.
+Because CLIP focuses more on task completion rather than comprehending the robotic manipulation process.
+LIV shows competence primarily in tasks involving obvious task goal definition.
+Our RoboCoT consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as drawers, windows, and doors.
+These results underscore the effectiveness of our RoboCoT in handling diverse and challenging scenarios.
+V-B Ablation Study
+In this section, we conduct ablation studies by removing each component of our algorithm including without self-imitation procedure, without sparse failure guidance and without robot Chain-of-Thought.
+The results indicate that each component of our RoboCoT contributes to the overall performance.
+In addition, we conduct ablation studies comparing GPT-4V with ground truth labels in collecting success experience, demonstrating the robustness of GPT-4V labeling.
+Finally, experiments with all baselines incorporating VLM-based self imitation procedure reveal that our algorithm’s reward design mechanism is superior compared to other baselines.
+Question 1: how significant is each component of our algorithm in contributing to the overall performance?
+Our algorithm incorporates several critical components: VLM-based self imitation procedure, sparse human failure guidance, and robot Chain-of-Thought process to define the rewards.
 To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance.
-A decline in performance upon the removal of any component demonstrates its essential role in our algorithm.
-Figure 4 illustrates the success rates of the final policy performance of our algorithm compared to its ablated versions.
-The labels for these ablations are: w/oReg (without success experience regularization, i.e., self imitation), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
-The results confirm the critical importance of each component in our algorithm, with VLM-based self imitation being particularly pivotal.
-Human failure guidance is the second most important component, motivating the agent to consider potential failure cases.
-Additionally, robot chain-of-thought processing helps the agent better understand the process of manipulation tasks, further contributing to the performance of our algorithm.
-Question 2: what is the empirical difference between the labels assigned by GPT-4V and the ground truth labels in the context of success experience collection?
-Figure 5: Ablation study for our algorithm with different success labeling methods (GPT-4V labeling or ground truth (GT) labeling).
-The results demonstrate that the success labels generated by GPT-4V are robust.
-Figure 6: Ablation study for all baselines with success experience regularization.
-The results show that our algorithm performs well not only because of the VLM-based self imitation component but also due to the reward design mechanism.
-Our algorithm RoboCoT incorporates GPT-4V to gather successful trajectories for VLM-based self imitation procedure.
-To validate the robustness of our GPT-4V labeling in collecting success experiences, we conduct experiments comparing GPT-4V labels and ground truth labels.
-If the performance using GPT-4V labels is similar to that using ground truth labels, it would confirm that GPT-4V is robust in collecting success experiences.
+A decline in performance upon the removal of any component demonstrates its essential role in our RoboCoT.
+Figure 5: Ablation study for our algorithm with different success experience collection methods (GPT-4V or ground truth (GT) labels).
+The results demonstrate that the GPT-4V are robust in collecting success experience in a subset of robotic tasks.
+Figure 6: Ablation study for all baselines with VLM-based self imitation procedure (w/reg). The results show that the reward design mechanism in our RoboCoT is superior than other baselines.
+Figure 4 illustrates the success rates of the final policy performance of our RoboCoT compared to its ablated versions.
+The labels for these ablations are without self-imitation procedure, without failure guidance and without Chain-of-Thought process.
+The results confirm the critical importance of each component in our algorithm.
+Each component in our RoboCoT contributes to the overall performance.
+In particular, VLM-based self imitation plays an important role in our RoboCoT.
+For better evaluating the importance of reward design mechanism in RoboCoT, we also conduct an ablation study with all baselines incorporating VLM-based self imitation procedure in Question 3.
+Question 2: Is GPT-4V robust in the context of success experience collection?
+Our algorithm RoboCoT incorporates GPT-4V to collect success experience for VLM-based self imitation procedure.
+To validate the robustness of our GPT-4V labels in collecting success experience, we conduct experiments comparing GPT-4V labels to ground truth labels.
 The process for GPT-4V analysis is outlined in Figure 2.
@@
 Otherwise, the output is 0." Subsequently, we present the final observation to GPT-4V and prompt, "What is the output of this image?" We collect the trajectory as successful when GPT-4V’s response is 1.
-Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labels for success experience collection.
-Our findings show that the GPT-4V labels achieves comparable performance to the ground truth across various tasks, demonstrating the robustness of our GPT-4V labels.
-Question 3: what is the empirical difference between our algorithm and all baselines with VLM-based self imitation?
-Question 1 shows that VLM-based self imitation component plays the most important role in our algorithm.
-To evaluate the significance of the reward design mechanism, we conduct experiments applying VLM-based self imitation procedure to all baselines.
-If the results show that our algorithm still outperforms other baselines with self imitation procedure, then it proves that our algorithm’s superiority is due not only to the VLM-based self imitation procedure but also to the reward design mechanism.
-Figure 6 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include VLM-based self imitation procedure.
+Figure 5 presents the training success rate curves comparing GPT-4V and ground truth (GT) labels for collecting success experience.
+Our findings indicate that GPT-4V’s labeling performance declines in part of the tasks.
+However, in most tasks, the performance is comparable to that of the ground truth labels, demonstrating that GPT-4V labels are robust for a subset of robotic tasks.
+Question 3: Is the reward design mechanism superior in our RoboCoT compared to other baselines?
+Question 1 shows that VLM-based self imitation component plays an important role in our RoboCoT.
+To evaluate the significance of the reward design mechanism, we conduct experiments incorporating VLM-based self imitation procedure to all baselines.
+If the results show that our algorithm still outperforms other baselines with VLM based self imitation procedure, then it proves that our algorithm’s reward design mechanism is superior than other baselines.
+Figure 6 shows the empirical results comparing our RoboCoT with other state-of-the-art baselines which also incorporating VLM-based self imitation procedure.
 The results demonstrate that our algorithm outperforms other baselines on almost all tasks.
-This indicates that our algorithm’s superiority is due not only to the VLM-based self imitation component but also to the reward design mechanism.
-6 Conclusion
-In this study, we present a novel algorithm RoboCoT, which defines rewards for skill acquisition.
-Our algorithm RoboCoT simplifies complex tasks through a robot’s chain-of-thought process with sparse failure guidance.
-It yields a more fine-grained reward guidance.
-Through VLM-based self imitation, our algorithm strengthens its policies, thereby improving performance.
-Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage over established baseline models.
-This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic tasks in the future.
-7 Limitation
-Our proposed method, RoboCoT, effectively masters a wide range of manipulation tasks through a robot chain-of-thought with sparse failure guidance.
-It provides a fine-grained reward guidance for the robot agent, facilitating the training of various automation tools.
-However, these tasks primarily focus on aligning robot movements with detailed language descriptions.
-RoboCoT does not yield promising results in tasks that emphasize observational preferences, such as soccer.
+It indicates that the reward design mechanism in our RoboCoT is superior in skill learning compared to other existing baselines.
+VI CONCLUSIONS
+In this study, we present a novel skill learning algorithm, i.e., RoboCoT, which provides more fine-grained reward guidance for learning manipulation skills.
+Our RoboCoT simplifies complex tasks through a robot chain-of-thought process.
+It motivates the robot to better comprehend the complex tasks for skill learning.
+Through VLM-based self imitation procedure, our RoboCoT speeds up the learning process, thereby improving its performance.
+Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage compared to other state-of-the-art baselines.
+This research sets a new standard for mastering manipulation skills in robotics.
+We hope our RoboCoT can be applied into more complex robotic tasks in the future.